Jan and Stewart Jones plan to borrow $20,000 for a new car. They are trying to decide whether to take out a 4-year or 5-year simple interest loan. The 4-year loan has an interest rate of 6% and the 5-year loan has an interest rate of 6.25%.
a. How much will they pay in interest on the 4-year loan?
A. $4500
B. $4800
C. $5000
D. $5200
b. How much will they repay with the 4-year loan?
A. $24,500
B. $24,800
C. $25,000
D. $25,200
c. How much more interest will they pay with the 5-year loan?
A. $5000
B. $6000
C. $6250
D. $6500
d. If the Jones can get a 5-year loan with 5.75% simple interest, which of the loans is the best deal? (lowest total cost)
A. $25,000
B. $26,000
C. $26,250
D. $26,500
e. How much more interest will they pay with the five-year loan?
A. $1000
B. $1450
C. $1500
D. $2000
f. If the Stewarts can get a 5-year loan with 5.75% simple interest, which of the loans is the best deal?
A. 4-year, 6%
B. 5-year, 5.75%
C. 5-year, 6.25
D. Cannot be determined
Thank you for dedicating some of your time to help me out with my homework. Showing your work is much appreciated. Have a nice day.

Answer by Boreal(15235)

Interest=principal*rate*time
a)I=20000*0.06*4=$4800
B
--------------------------
b)Repay $24,800.
B
--------------------------
c)
I=20,000*5*0.0625=$6250. How much more interest will they pay? Relative to the 4 year loan, which had interest of $4800, they are paying $1450 more, which isn't one of the choices.
------------
d)20000*0.575*5=$5750
This isn't one of the choices.
------------
-------------------------
e)5 year loan is $6250 interest-$4800 with 4 year=$1450 difference
B
------------------------
f)I=20000*5*0.0575=$5750
Repay $25,750
This is a smaller rate than the 4 year loan, but it is spread out over an additional year, where the money could not be invested in something else. Cannot be determined.
D